# thanks: ORCID ID: 0009-0003-9679-035Xthanks: ORCID ID: $0$-0001-7989-9271

### v1

@@ 3 Self-propulsion mechanism
Figure 3: Properties of the local reactor (11). Panels (a) and (b) show the phase portraits outside and inside the Hopf unstable region with $\mu_{2}=3.5$ and $\mu_{2}=4.8$. The nullcline of $\Gamma_{1}$ ($\Gamma_{2}$) is represented as a blue (green) line. The stable (unstable) fixed point is shown as a filled (empty) point. Red lines represent typical trajectories, in (a) the red square denotes the initial condition, in (b) the trajectory corresponds to the limit cycle. Panels (c) and (d) show the time evolution of $\Gamma_{1}$ (blue) and $\Gamma_{2}$ (green) for the red trajectories in (a) and (b), respectively. In (c), the squares denote the initial concentrations. Panel (e) shows the numerically computed bifurcation diagram with $\beta_{2}e^{\mu_{2}}$ as control parameter ($\beta_{2}$ fixed). The steady state concentrations $\Gamma_{1,ss}$ and $\Gamma_{2,ss}$ are drawn in blue and green (dotted lines) the sum of both concentrations is shown in orange [black for estimate given by Eqs. (12)]. Supercritical Hopf bifurcations occur at the extrema of $\Gamma_{1,ss}+\Gamma_{2,ss}$ (orange dots), between which the steady state is unstable (dashed line). In the unstable region, a limit cycle exists, where the mean total concentration is represented as a purple solid line. The parameter values for (a) and (b) are marked. (f) Dependence of the steady state surface tension $\gamma(\Gamma_{1,ss}+\Gamma_{2,ss})$ on both driving currents $\beta_{1}e^{\mu_{1}},\beta_{2}e^{\mu_{2}}$ (fixed $\beta_{1},\beta_{2}$) given by Eqs. (12). The stability regions shown at the bottom are given by Eq. (13). In Panels (e) and (f), pentagon [triangle] markers denote the parameter choices in (a) and (c) [(b) and (d)]. The remaining parameters for all panels are $r=0.3,\beta_{1}=2,\beta_{2}=0.01,\delta=1$ and [except (e)] $\mu_{1}=-1.4$.
@@
First, we consider 2D droplets (liquid ridges) on a one-dimensional domain with periodic boundary conditions, and study the dynamics using finite-element based time simulations implemented in oomph-lib [65] (Appendix E). We specifically examine scenarios of strong molecular interactions between the liquid and the substrate as compared to the energetic influence of surfactant ($W=\frac{Aa_{1}a_{2}}{h_{a}^{2}k_{b}T}\gg 1$, Appendix B). In both the passive and the active case, flat films then typically rupture by spinodal dewetting (i.e., by a long-wave instability of Cahn-Hilliard type, see Supplemental Material of [66] for a recent classification) and quickly form a single droplet [Fig. 2(b) and Fig. 2(c)]. At large driving forces also other instability types can occur [Fig. 2(c)], although here we focus on the spinodal scenario. If the ambient bath is depleted of the autocatalyst $\Gamma_{1}$ and rich in the reactant $\Gamma_{2}$ ($\mu_{1}<0,\mu_{2}>0$), droplets formed by dewetting spontaneously break their left-right symmetry and move across the substrate [Fig. 2(a) and Fig. 2(b), Supplemental Video 1]. This is due to Marangoni convection induced by a greatly increased surface tension in the advancing contact line region. The surface tension in the receding contact line region is also slightly increased. The effect at the receding contact line vanishes when the coupling of the liquid pressure to the surfactant dynamics [see Eq. (8)] is neglected and results from pressure gradients in the contact line regions. Because this effect is comparatively small we do not discuss it further.
@@
Importantly, self-propelled droplets can be observed even when both surfactants diffuse equally ($D_{1}\!=\!D_{2}$) which excludes a Turing instability of the reaction-diffusion subsystem as an underlying mechanism. We now show that the mechanism for generating and maintaining local gradients in surface tension arises from the chemo-mechanical interplay of the nonlinear “local reactor” and Marangoni convection. Ultimately, this gives rise to various forms of self-propelled droplets. To this end, temporarily we only consider the (nondimensional) local reactor given by
@@
where $\delta=a_{1}/a_{2}$. To compute the steady states of the system given by (11), we additionally assume that $\beta_{2}=\mathcal{O}(\varepsilon)$ and $e^{\mu_{2}}=\mathcal{O}(\varepsilon^{-1})$ with $\varepsilon\ll 1$ and that all other quantities are $\mathcal{O}(1)$. This reflects a slow exchange of $\Gamma_{2}$ with the bath and a strong energetic bias towards the adsorption of $\Gamma_{2}$ onto the droplet. These assumptions also capture the parameter choice of Fig. 2. Then, the local reactor (11) has a single fixed point which can be determined to order $\mathcal{O}(1)$ as
@@
$\frac{\partial\left(\Gamma_{1,ss}+\Gamma_{2,ss}\right)}{\partial(\beta_{2}e^{\mu_{2}})}=0.$ (13)
@@
That is, the Hopf bifurcations occur at the local extrema of $\gamma_{ss}\left(\beta_{2}e^{\mu_{2}}\right)$ [see Fig. 3(e)]. [^2] This condition can be understood intuitively by considering a two-component dynamical system with an N-shaped (e.g., cubic) and a vertical nullcline. This is the approximate shape of the nullclines of Eqs. (11) near the fixed point. In this case, the Hopf bifurcations generically occur when the fixed point crosses the extrema of the N-shaped nullcline. Approximately at these points, also $\Gamma_{1}+\Gamma_{2}$ exhibits a local extremum along the branch of steady states that is obtained by shifting the two nullclines with respect to each other. Inbetween, chemical oscillations are observed [Figs. 3(b) and 3(d)]. Outside of this parameter region, all phase space trajectories converge to the fixed point [Fig. 3(a), Fig. 3(c)]. We find that the corresponding reaction-diffusion system shows similar dynamics when $D_{1}=D_{2}$ and relaxes to the homogeneous steady state if it is stable.
@@
Figure 4: (a) Magnification of the front of a self-propelled droplet that moves with velocity $v$. The top panel shows the film thickness $h$ and streamlines represent the velocity field of the liquid in the laboratory frame. The bottom panel shows the surface tension profile $\gamma$. The three regions correspond to (i) the bulk of the droplet, (ii) the advancing contact line and (iii) the adsorption layer. In (i) and (iii), the surface tension $\gamma_{1}$ is given by the steady state of the local reactor, i.e., by evaluating $\gamma_{ss}$ at $\beta_{2}e^{\mu_{2}}$. At the surface tension maximum $z=z_{m}$ in (ii), the surface tension $\gamma_{2}$ is determined as a balance between transport fluxes and chemical reactions and can be obtained by evaluating $\gamma_{ss}$ at $\beta_{2}e^{\mu_{2}}\!-\!\partial_{z}j_{2}(z_{m})$. This is illustrated in (b), where $\gamma_{ss}$ is shown as a function of $\beta_{2}e^{\mu_{2}}$. The gray region marks the section between the extrema of $\gamma_{ss}$ (orange dots) where the steady state surface tension increases with increasing influx $-\partial_{z}j_{2}$. Note that $\gamma_{1}$ lies outside of this region. This capacity of the local reactor to increase the surface tension in response to an influx of reactant suggests a general positive feedback loop that is shown in (c).
@@
We now turn back to the complete spatially extended system and consider a self-propelled droplet in the comoving frame $z=x-vt$. If the droplet moves with constant velocity $v$, the surfactant profiles are given by
00 0 $=-\partial_{z}j_{1}+\mathcal{R}+\mathcal{B}_{1}-v\partial_{z}\Gamma_{1},$ (14)
00 0 $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2}-v\partial_{z}\Gamma_{2}.$
@@
We consider self-propelled states given by Eqs. (14) in three different regions (i) to (iii) shown in Fig. 4(a). Region (i) is located away from the contact line region and corresponds to the bulk of the droplet. The flow is nearly laminar and matter is “passed through” to the contact line region. Consequently, no surfactant accumulates due to transport and we have $\partial_{z}j_{1}=\partial_{z}j_{2}=0$. We find that the surfactant profile is uniform in this region such that $v\partial_{z}\Gamma_{1}=v\partial_{z}\Gamma_{2}=0$. The concentrations are then given by the steady state of the local reactor (12) and we denote the respective surface tension by $\gamma_{1}$. Region (iii) comprises the adsorption layer far away from the contact line region. Here, the film thickness is constant ($h=1$, or $h=h_{a}$ in dimensional units) and the surfactant coverage is uniform. As before, we then have $\partial_{z}j_{i}=v\partial_{z}\Gamma_{i}=0$ with $i=1,2$ and the local concentrations are again given by Eqs. (12) with the surface tension $\gamma_{1}$. Finally, we turn to the contact line region (ii). There, the liquid flow is dominated by two vortices that are “squeezed” into the contact line region due to the strong left-right symmetry breaking. The weaker vortex reaches into the adsorption layer [Fig. 4(a)]. Here, the fluxes $j_{1}$ and $j_{2}$ are generally not constant and there exist strong surface tension gradients. However, we only consider the local maximum of the surface tension profile at $z=z_{m}$ with the surface tension $\gamma_{2}$ and where $\partial_{z}\gamma=0$ and approximately $\partial_{z}\Gamma_{1}=\partial_{z}\Gamma_{2}=0$ (since the local extrema of $\Gamma_{1}$ and $\Gamma_{2}$ do not coincide exactly). In particular when the droplet speed is not too large, the contributions $-v\partial_{z}\Gamma_{1},-v\partial_{z}\Gamma_{2}$ are then negligible. We additionally observe in time simulations that the transport contributions $-\partial_{z}j_{1}$ are small near the surface tension peak. This can be explained by observing that $\Gamma_{1}$ is enriched in the contact line region [see Fig. 2(a)] such that diffusion of $\Gamma_{1}$ therefore acts in opposition to the advective fluxes (which transport surfactant into the contact line region). Diffusive and advective contributions then effectively cancel near the peak. We further discuss these approximations in Appendix D. As a result, at the local maximum of the surface tension profile at $z=z_{m}$ we have the balance equations
00 0 $=\mathcal{R}+\mathcal{B}_{1},$ (15)
00 0 $=-\mathcal{R}+\mathcal{B}_{2}-\partial_{z}j_{2}.$
@@
We recognize that Eqs. (15) correspond to the steady state equations of the local reactor (11), augmented by the spatial transport of reactant $\Gamma_{2}$. We are now interested in steady states with $-\partial_{z}j_{2}(z_{m})>0$ at the peak while $\Delta\gamma=\gamma_{2}-\gamma_{1}>0$. In this scenario matter is continuously advected into the contact line region due to a simultaneously maintained gradient in surface tension between the three discussed regions, which must hold for self-propelled droplets. Using Eqs. (15), we see that $-\partial_{z}j_{2}(z_{m})$ (which can be directly determined from time simulations) acts simply as an additional driving term that may be added to the control parameter, i.e., at the surface tension peak the surface tension $\gamma_{2}$ can be found by evaluating $\gamma_{ss}$ as a function of $\beta_{2}e^{\mu_{2}}\!-\!\partial_{z}j_{2}$. Because $\gamma_{ss}$ changes non-monotonically with the control parameter, there is a region of the curve where $\gamma_{ss}$ increases with the driving current, i.e., the total mass in the reactor decreases with an increasing influx of reactant [Fig. 4(b)]. We then find that $\Delta\gamma=\gamma_{ss}\left(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m})\right)-\gamma_{ss}\left(\beta_{2}e^{\mu_{2}}\right)>0$ may coincide with $-\partial_{z}j_{2}(z_{m})>0$ for sufficiently large transport contributions (note that the stability of the local reactor does not reflect the stability of the self-propelled state). Physically, this effect can be summarized as follows. The local reactor may overcompensate additional influxes of the reactant $\Gamma_{2}$ from neighboring regions on the free surface by a net removal of surfactant from the droplet, thereby maintaining a locally increased surface tension. In consequence, matter is continuously advected into the advancing contact line region and the droplet moves across the substrate. Self-propelled droplets as shown in Fig. 2 and Fig. 4 therefore represent a balance between chemical reactions and transport fluxes that is sustained due to an interplay of the nonlinear local reactor and the Marangoni effect.

### v3

@@ 3 Self-propulsion mechanism
Figure 3: Properties of the local reactor (11). Panels (a) and (b) show the phase portraits outside and inside the Hopf unstable region with $\mu_{2}=3.3$ and $\mu_{2}=4.8$, respectively. The nullcline of $\Gamma_{1}$ ($\Gamma_{2}$) is represented as a blue (green) line. The stable (unstable) fixed point is shown as a filled (empty) circle symbol. Red lines represent typical trajectories, in (a) the red square denotes the initial condition, in (b) the trajectory corresponds to the limit cycle. Panels (c) and (d) show the time evolution of $\Gamma_{1}$ (blue) and $\Gamma_{2}$ (green) for the red trajectories in (a) and (b), respectively. In (c), the squares denote the initial concentrations. Panel (e) shows the numerically computed bifurcation diagram with $\beta_{2}e^{\mu_{2}}$ as control parameter ($\beta_{2}$ fixed). The steady state concentrations $\Gamma_{1,ss}$ and $\Gamma_{2,ss}$ are drawn in blue and green (dotted lines) the sum of both concentrations is shown in orange [black for estimate given by Eqs. (12)]. Supercritical Hopf bifurcations occur at the extrema of $\Gamma_{1,ss}+\Gamma_{2,ss}$ (orange dots). Between them, the steady state is unstable (dashed line). In the unstable region, a limit cycle exists, where the mean total concentration is represented as a purple solid line. (f) Dependence of the steady state surface tension $\gamma(\Gamma_{1,ss}+\Gamma_{2,ss})$ on both driving currents $\beta_{1}e^{\mu_{1}},\beta_{2}e^{\mu_{2}}$ (fixed $\beta_{1},\beta_{2}$) given by Eqs. (4) and (12). The stability diagram shown at the bottom results from Eq. (13). In Panels (e) and (f), pentagon [triangle] markers denote the parameter choices in (a) and (c) [(b) and (d)]. The remaining parameters for all panels are $r=0.3,\beta_{1}=2,\beta_{2}=0.01,\delta=1$ and [except (f)] $\mu_{1}=-1.4$.
@@
First, we consider 2D droplets (liquid ridges) on a one-dimensional domain with periodic boundary conditions, and study the dynamics using finite-element based time simulations implemented in oomph-lib [75] (Appendix E). We specifically examine scenarios of strong molecular interactions between the liquid and the substrate as compared to the energetic influence of surfactant ($W=\frac{Aa_{1}a_{2}}{h_{a}^{2}k_{b}T}\gg 1$, Appendix B). In both the passive and the active case, flat films then typically rupture by spinodal dewetting (i.e., by a long-wave instability of Cahn-Hilliard type, see Supplemental Material of [76] for a recent classification) and quickly form a single droplet [Fig. 2(b) and Fig. 2(c)]. At large driving force also other instability types can occur [Fig. 2(c)], although here we focus on the spinodal scenario. If the ambient reservoir is depleted of the autocatalyst $\Gamma_{1}$ and rich in the reactant $\Gamma_{2}$ ($\mu_{1}<0,\mu_{2}>0$), droplets formed by dewetting spontaneously break their left-right symmetry and move across the substrate [Fig. 2(a) and Fig. 2(b), Supplemental Video 1]. This is due to Marangoni convection induced by a greatly increased surface tension in the advancing contact line region. The surface tension in the receding contact line region is also slightly increased. The effect at the receding contact line vanishes when the coupling of the liquid pressure to the surfactant dynamics [see Eq. (8)] is neglected and results from pressure gradients in the contact line regions. Because this effect is comparatively small we do not discuss it further.
@@
Importantly, self-propelled droplets can be observed even when both surfactants diffuse equally ($D_{1}\!=\!D_{2}$) which excludes a Turing instability of the reaction-diffusion subsystem as an underlying mechanism. We now show that the mechanism for generating and maintaining local gradients in surface tension arises from the chemomechanical interplay of the nonlinear “local reactor” and Marangoni convection. Ultimately, this gives rise to various forms of self-propelled droplets. To this end, temporarily we only consider the (nondimensional) local reactor given by
@@
where the dot denotes the derivative with respect to time and $\delta=a_{1}/a_{2}$. To compute the steady states of the system given by (11), we additionally assume that $\beta_{2}=\mathcal{O}(\varepsilon)$ and $e^{\mu_{2}}=\mathcal{O}(\varepsilon^{-1})$ with $\varepsilon\ll 1$ and that all other quantities are $\mathcal{O}(1)$. This reflects a slow exchange of $\Gamma_{2}$ with the bath and a strong energetic bias towards the adsorption of $\Gamma_{2}$ onto the droplet. These assumptions also capture the parameter choice of Fig. 2. Then, the local reactor (11) has a single fixed point which can be determined to order $\mathcal{O}(1)$ as
@@
$\frac{\partial\left(\Gamma_{1,ss}+\Gamma_{2,ss}\right)}{\partial(\beta_{2}e^{\mu_{2}})}=0,$ (13)
@@
where both $\Gamma_{1,ss}$ and $\Gamma_{2,ss}$ are understood as functions of $\beta_{2}e^{\mu_{2}}$ and of all other parameters (except $\beta_{2}$ and $\mu_{2}$) appearing in Eqs. (12). That is, the Hopf bifurcations occur at the local extrema of $\gamma_{ss}\left(\beta_{2}e^{\mu_{2}}\right)$ [see Fig. 3(e)]. [^2] This condition can be understood intuitively by considering a two-component dynamical system with an N-shaped (e.g., cubic) and a vertical nullcline. This is the approximate shape of the nullclines of Eqs. (11) near the fixed point. In this case, the Hopf bifurcations generically occur when the fixed point crosses the extrema of the N-shaped nullcline. Approximately at these points, also $\Gamma_{1}+\Gamma_{2}$ exhibits a local extremum along the branch of steady states that is obtained by shifting the two nullclines with respect to each other. In between, chemical oscillations are observed [Figs. 3(b) and 3(d)]. Outside of this parameter region, all phase space trajectories converge to the fixed point [Fig. 3(a), Fig. 3(c)]. We find that the corresponding reaction-diffusion system shows similar dynamics when $D_{1}=D_{2}$ and relaxes to the homogeneous steady state if it is stable. Note that the Hopf bifurcations given by condition (13) correspond exactly to the Hopf bifurcations of the flat film (Fig. 2(c) shows the first). Therefore, due to our restriction to the purely spinodal case shown in Fig. 2(c), we are operating within this stable regime of the local reactor also for the full model (7).
@@
Figure 4: (a) Magnification of the front of a self-propelled droplet that moves with velocity $v$. The top panel shows the film thickness $h$ and streamlines represent the velocity field of the liquid in the laboratory frame. The bottom panel shows the surface tension profile $\gamma$. The three regions correspond to (i) the bulk of the droplet, (ii) the advancing contact line and (iii) the adsorption layer. In (i) and (iii), the surface tension $\gamma_{1}$ is given by the steady state of the local reactor, i.e., by evaluating $\gamma_{ss}$ at $\beta_{2}e^{\mu_{2}}$. At the surface tension maximum $z=z_{m}$ in (ii), the surface tension $\gamma_{2}$ is determined as a balance between transport fluxes and chemical reactions and can be obtained by evaluating $\gamma_{ss}$ at $\beta_{2}e^{\mu_{2}}\!-\!\partial_{z}j_{2}(z_{m})$. This is illustrated in (b), where $\gamma_{ss}$ is shown as a function of $\beta_{2}e^{\mu_{2}}$. The gray region marks the section between the extrema of $\gamma_{ss}$ (orange dots) where the steady state surface tension increases with increasing influx $\partial_{z}j_{2}$. Note that $\gamma_{1}$ lies outside of this region. This capacity of the local reactor to increase the surface tension in response to an influx of reactant suggests a general positive feedback loop that is shown in (c).
@@
We now turn back to the complete spatially extended system and consider a self-propelled droplet in the comoving frame z=xvtz=x-vt. If the droplet moves with constant velocity $v$, the surfactant profiles are given by
$0$ $=-\partial_{z}j_{1}+\mathcal{R}+\mathcal{B}_{1}-v\partial_{z}\Gamma_{1},$ (14)
$0$ $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2}-v\partial_{z}\Gamma_{2}.$
@@
We consider self-propelled states given by Eqs. (14) in three different regions (i) to (iii) shown in Fig. 4(a). Region (i) is located away from the contact line region and corresponds to the bulk of the droplet. The flow is nearly laminar and matter is “passed through” to the contact line region. Consequently, no surfactant accumulates due to transport and we have $\partial_{z}j_{1}=\partial_{z}j_{2}=0$. We find that the surfactant profile is uniform in this region such that $v\partial_{z}\Gamma_{1}=v\partial_{z}\Gamma_{2}=0$. The concentrations are then given by the steady state of the local reactor (12) and we denote the respective surface tension by $\gamma_{1}$. Region (iii) comprises the adsorption layer far away from the contact line region. Here, the film thickness is constant ($h=1$, or $h=h_{a}$ in dimensional units) and the surfactant coverage is uniform. As before, we then have $\partial_{z}j_{i}=v\partial_{z}\Gamma_{i}=0$ with $i=1,2$ and the local concentrations are again given by Eqs. (12) with the surface tension $\gamma_{1}$. Finally, we turn to the contact line region (ii). There, the liquid flow is dominated by two vortices that are “squeezed” into the contact line region due to the strong left-right symmetry breaking. The weaker vortex reaches into the adsorption layer [Fig. 4(a)]. Here, the fluxes $j_{1}$ and $j_{2}$ are generally not constant and there exist strong surface tension gradients. However, we only consider the local maximum of the surface tension profile at $z=z_{m}$ with the surface tension $\gamma_{2}$ and where $\partial_{z}\gamma=0$ and approximately $\partial_{z}\Gamma_{1}=\partial_{z}\Gamma_{2}=0$ (since the local extrema of $\Gamma_{1}$ and $\Gamma_{2}$ do not coincide exactly). In particular when the droplet speed is not too large, the contributions $v\partial_{z}\Gamma_{1},-v\partial_{z}\Gamma_{2}$ are then negligible. We additionally observe in time simulations that the transport contributions $\partial_{z}j_{1}$ are small near the surface tension peak. This can be explained by observing that $\Gamma_{1}$ is enriched in the contact line region [see Fig. 2(a)] such that diffusion of $\Gamma_{1}$ opposes the advective fluxes (which transport surfactant into the contact line region). Diffusive and advective contributions then effectively cancel near the peak. Note that the difference in coverage between $\Gamma_{1}$ and $\Gamma_{2}$ (particularly near the advancing contact line) originates from differences in $\beta_{1},\beta_{2}$ and $\mu_{1},\mu_{2}$ as well as from the distinct roles of the two species in the autocatalysis (5) since we assume $D_{1}=D_{2}$ and $\delta=1$. We further discuss these approximations in Appendix D. As a result, at the local maximum of the surface tension profile at $z=z_{m}$ we have the balance equations
$0$ $=\mathcal{R}+\mathcal{B}_{1},$ (15)
$0$ $=-\mathcal{R}+\mathcal{B}_{2}-\partial_{z}j_{2}.$
@@
We recognize that Eqs. (15) correspond to the steady state equations of the local reactor (11), augmented by the spatial transport of reactant $\Gamma_{2}$. We are now interested in steady states with $\partial_{z}j_{2}(z_{m})>0$ at the peak while $\Delta\gamma=\gamma_{2}-\gamma_{1}>0$. In this scenario matter is continuously advected into the contact line region due to a simultaneously maintained gradient in surface tension between the three discussed regions, which must hold for self-propelled droplets. Using Eqs. (15), we see that $\partial_{z}j_{2}(z_{m})$ (which can be directly determined from time simulations) acts simply as an additional driving term that may be added to the control parameter, i.e., at the surface tension peak the surface tension $\gamma_{2}$ can be found by evaluating $\gamma_{ss}$ as a function of $\beta_{2}e^{\mu_{2}}\!-\!\partial_{z}j_{2}$. Because $\gamma_{ss}$ changes non-monotonically with the control parameter, there is a region of the curve where $\gamma_{ss}$ increases with the driving current, i.e., the total mass in the reactor decreases with an increasing influx of reactant [Fig. 4(b)]. We then find that $\Delta\gamma=\gamma_{ss}\left(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m})\right)-\gamma_{ss}\left(\beta_{2}e^{\mu_{2}}\right)>0$ may coincide with $\partial_{z}j_{2}(z_{m})>0$ for sufficiently large transport contributions (note that the stability of the local reactor does not reflect the stability of the self-propelled state). Physically, this effect can be summarized as follows. The local reactor may overcompensate additional influxes of the reactant $\Gamma_{2}$ from neighboring regions on the free surface by a net removal of surfactant from the droplet, thereby maintaining a locally increased surface tension. In consequence, matter is continuously advected into the advancing contact line region and the droplet moves across the substrate. Self-propelled droplets as shown in Fig. 2 and Fig. 4 therefore represent a balance between chemical reactions and transport fluxes that is sustained due to an interplay of the nonlinear local reactor and the Marangoni effect.